### v3

@@ 4.1 Optical model and simulation
The interaction and propagation of particles in the scintillator volume were simulated with the Geant4 toolkit [79, 80, 81]. The polyvinyltoluene-based plastic scintillator is simulated with light yield (8,700 photons / MeV), emission and absorption spectra, attenuation length (2.5 m), refractive index (1.58), and decay time ( ns) characteristic of EJ-262 [66], the scintillator used for the PLATON prototype. Initial-state particles can be either generated using Geant4 as single particles or, for neutrinos, using the NEUT generator [71]. We assume the distance between the position of the ionising particle and the final emission of the scintillation light to be small compared to the 3D spatial resolution of the PLATON camera [82, 16] and, thus, no additional smearing on the position of the scintillation light emission is applied. The optical photons produced by the scintillation process are propagated through the scintillator volume until the edge using the native Geant4 optical photon processes, and their positions and angles are stored upon exiting the scintillation volume.
The second part of the optical simulation is based on a custom library that simulates lenses, sensors, and refractive surfaces outside the scintillator volume. It performs the ray tracing of each individual photon from the exit point on the scintillator surface to the photosensor, through the plenoptic system. Thin or thick lenses, along with their aperture, are implemented in the paraxial approximation. General surfaces can be added to the optical stack by defining their shape using tessellated triangles in an STL file.
@@
The simulation of the SPAD array sensors was based on the parameters of SwissSPAD2 [83]. The pixel active and dead areas of the SPAD array were simulated as designed, thus with a geometrical fill factor. The photon detection efficiency (PDP) was simulated as a binomial efficiency following the corresponding wavelength spectrum. The DCR was also cross-checked against measurements performed in our laboratory, and where explicitly mentioned, extrapolated to a future, realistic performance. The timestamp of each pixel can be simulated in either self-trigger or gate mode.
The positions and the relative distances of the optical objects were adjusted to reflect the configuration of the PLATON prototype (see Sec. 4.2 and 4.3) or that of an optimised PLATON detector (see Sec. 2.3). Although this model appears rather simple, it accurately reproduces the actual resolution of the PLATON prototype.
@@ 4.2 Image post-processing method
Post-processing algorithms for plenoptic systems are based on parallax detection and require pattern matching between images observed by different micro-lenses [84]. However, we decided not to adopt this technique as it cannot be easily applied to photon-starved frames, as it would lead to the pattern matching failing. Thus, we require an optical model (see Sec. 4.1) to project each detected photon through the optics into the object space. Post-processing, based on that of a pinhole camera, is also used for plenoptic cameras in light-intensity imaging [85]. In this work, we adopt the same principle to the case of photon-starved images and, thus, developed a new method: the ray trace of every photon that was detected by a pixel of the sensor (a count) is forced to pass through the centre of the closest micro-lens, and interpolated in the object space. The optical model described in Sec. 4.1 was used for the propagation of the ray tracing: the MLA was parametrised as an array of pinholes, and each photon was assumed to have impinged at the centre of the pixel. Outlier ray traces, mainly generated by noise, but also by the adopted approximation, were iteratively rejected until either a predefined minimum number of rays with the relative closest distance was obtained or the distance between each leftover ray was below a certain threshold. The interpolation with the remaining ray traces was used to estimate the position of the light source.
@@
4.3 Calibration of the PLATON prototype
The post-processing of the 2D images relies on the accurate calibration of the optical model of the prototype, that is, the data-driven inference of the model parameters discussed in Sec. 4.1. First, the lateral position of each micro-lens is obtained by illuminating the PLATON prototype without the main lens with a source of parallel light rays. This was achieved using a low-power laser mounted on a custom-made motorised stage, which could be translated parallel to the MLA plane. A single greyscale intensity image was obtained from 4’096 frames and used to determine the centre of each micro-lens. The mean distance between the centre of two adjacent micro-lenses was computed to derive the parameter . The position of the micro-lenses could not be determined for the entire area of the MLA, as the glue partially covers that region (see Fig. 1). Consequently, that part of the image was not used in the post-processing. This issue reduced the field of view of the PLATON camera. Imaging a parallel-light source is not sufficient to infer the remaining parameters of the optical model. Hence, we acquired a dataset consisting of 12-bit images of a back light-illuminated 50 m pinhole at 168 different positions within an imaging volume of . Each image was obtained by stacking 4,096 frames, each recorded with a 10 s gate. The motorised stage, by design, allowed the light source to move in steps of . Measurements where the light source fell outside the camera’s field of view (FoV), reduced by the glue, were removed from the dataset. In total, 160 images were used to fit the parameters of the optical model using a custom genetic algorithm, in which the post-processing method was iteratively applied and the distance between the reconstructed and true positions was minimised.
@@ 4.3 Calibration of the PLATON prototype
The optical simulation (see Sec. 4.1) was used to reproduce images from a 100 point source at the same 168 positions as those of the acquired dataset. The focal length of the microlens array, , was fixed to its nominal value in the datasheet of the MLA. Then, the simulated images were compared to the corresponding data samples, and good agreement was observed, indicating that the model provides a reasonable representation of the camera’s optical stack.
@@ 4.4 Data-driven likelihood reconstruction analysis
Alternative post-processing methods were also tested. The first one consists of a data-driven likelihood analysis. The intensity greyscale images for each of the 160 points scanned during the calibration campaign were used as probability density functions (PDFs). To each PDF, the distribution of the dark counts across the pixels, obtained with an independent measurement, is subtracted. This way we could estimate the intrinsic spatial resolution of the plenoptic system for the case of a negligible contamination from dark counts. This study aims to fake the scenario of a future SPAD array chip capable of rejecting dark counts by time coincidence. See more details in the discussion in Sec. 3. “Sampled” data frames were obtained by randomly throwing the position of the counts from the PDF of a given position of the light source.
@@
where the first term contributes only when the given pixel in the sampled frame takes the value of 1, while the second term contributes when the pixel is 0 and weighs the probability of the pixel to be off. This likelihood definition allows for taking into account both activated and non-activated pixels. For each one, the was computed with respect to the PDF at every scanned position , and the position yielding the lowest value was selected as the best-fit estimate. We refer to this approach as the “likelihood” method.
Since the testable resolution of the likelihood method depends on the distance between two adjacent scanned positions, the calibration sample was augmented within a smaller portion of the volume, which was sampled with a higher density of points. At the depths of 245 mm and 345 mm from the sensor, 101 16-bit images were taken respectively, with a spacing of 0.1 mm. These were used to determine the lateral resolution. For the depth resolution, data points with a spacing of 1 mm along the optical axis, between a 230 mm and 450 mm distance from the sensor, were gathered. Each of these intensity images was used as a PDF to infer the position of the point light source. The fake datasets were generated as described above.
@@ 4.4.1 Point-light spatial resolution and comparison with classical cameras
Before explicitly studying the capability of a PLATON module to detect GeV neutrinos, a simulation study was performed to evaluate the spatial resolution to a point light source corresponding to 1 MeV energy deposition. The post-processing method used to evaluate the spatial resolution for a point light source is the same as described in Sec. 4.2. For a single-point scintillation light source, creating 10,000 photons, the 3D spatial resolution is 0.3 mm. If the number of cameras is reduced to two, the 3D spatial resolution becomes 0.8 mm.
For consistency, we compared the performance of the plenoptic cameras with classical cameras, obtained by removing the MLA and adapting the lens-to-sensor distance. The post-processing adopted for the classical cameras is the same as that used for plenoptic cameras, with the difference that now the pinhole is the main lens. While a single classical camera cannot precisely reconstruct the 3D position of the light source, the resolution of two orthogonal cameras is 2.4 mm, which is three times worse than that of the corresponding plenoptic system. Moreover, we note that two classical cameras fail to reconstruct the position of the light source for approximately 23% of the events, whereas all events are successfully reconstructed for two plenoptic cameras. The resolution for eight cameras, split into two orthogonal views, is approximately four times worse than that of the PLATON detector module, which utilises eight plenoptic cameras.
@@ 4.5 Neural-network based reconstruction
The dataset used for training the neural network comprised CC-inclusive muon neutrino interactions generated as described in Sec. 2.3.2. The simulation of the detector setup is discussed in Sec. 2.3. Reconstructing 3D neutrino interactions within this setup is inherently challenging. SPAD arrays are arranged orthogonally on two sides of the scintillator block. Each interaction produces a photon-starved pattern, typically comprising between a few tens and several thousand detections across the SPADs. The task is to infer the spatial origin and topology of the interaction using only these sparse observations.
@@
To achieve this, we opt for a transformer-encoder-based architecture to fully exploit the spatial correlations among detected photons through the transformer’s attention mechanism [86, 87]. Unlike traditional architectures that rely on local connectivity or fixed sequence order, the transformer’s attention mechanism enables dynamic interactions between all elements in the input. This is particularly beneficial for our application, where the input consists of sparse and spatially-distributed photon detections and where meaningful correlations may exist between distant regions of the detector. Compared to convolutional neural networks (CNNs) [88, 89] or graph neural nets (GNNs) [90, 91, 92], which are constrained by local or explicitly defined receptive fields, transformers are not limited to such neighbourhoods and can integrate global spatial context more effectively. Recurrent neural networks (RNNs) [93, 94, 95] are also ill-suited for this task, as they assume a consistent sequential ordering, which does not naturally apply to unordered 2D photon hits [96]. In contrast, the transformer’s attention mechanism provides the flexibility to learn arbitrary spatial dependencies, which is essential for resolving the complex and often ambiguous photon distributions resulting from neutrino interactions [97, 98]. An alternative class of models worth considering is Neural Radiance Fields (NeRFs) [99], which learn continuous volumetric scene representations from dense multi-view RGB data using ray-based modelling. While NeRFs have proven effective for reconstructing 3D geometry from 2D projections, their direct application to our setting is limited. SPAD sensors produce extremely sparse, binary detections, lacking the dense photometric supervision and calibrated viewing geometries that NeRFs rely on. Furthermore, NeRFs are primarily designed for photorealistic view synthesis, whereas our goal is to recover the underlying spatial topology of particle interactions from non-image-based, unordered data. As such, although NeRF-inspired techniques may hold promise for future work [100, 101], they would require substantial re-engineering to address the unique challenges of our domain. Moreover, more traditional reconstruction methods are discarded due to the multi-dimensional complexity of the problem and the large number of sparse photons involved. These methods often lack the flexibility and computational efficiency required to process and interpret the intricate patterns of photon detection necessary for accurate 3D reconstruction.
@@
We employ a transformer encoder to tackle this reconstruction problem (see Figure 9)a. In our setup, each photon is treated as a token, embedding its 2D position on the SPAD array, the identifier of the microlens it passed through, and the corresponding SPAD sensor ID. Additionally, the input sequence includes a dedicated token encoding the number of sampled photons in the event and one token for each muon exiting the scintillator block (typically one), which contains the 3D exit point and direction vector. The maximum sequence length is fixed at 1,024 tokens. For events where the total number of tokens (including the muon and size tokens) would exceed this limit, we randomly sample photons to stay within the bound. Conversely, if the event has fewer photons, we randomly select 90% of them as a form of regularisation and fill the remainder of the sequence with a padding token, which is ignored during training. Given that each token already includes spatial information, explicit positional encodings are unnecessary. Early experiments showed that the network struggled to reconstruct events when the number of detected photons was significantly lower than the maximum sequence length. We hypothesise that this limitation arises from the model’s lack of awareness of the photon count in each event. To address this, we include the aforementioned dedicated token encoding the sequence size, which resolves this issue by enabling the model to condition the number of observed photons.
@@
The quadratic complexity of the standard transformer’s attention mechanism poses computational challenges for processing events with a large number of photons. To overcome this, we use a custom-adapted version of BigBird’s sparse attention [102], which combines global, window, and random attention within blocks of tokens to achieve linear complexity while effectively modelling long-range dependencies. Global attention allows specific tokens to attend to all others, window attention limits tokens to attend only to their neighbours, and random attention randomly connects tokens. This hybrid approach allows the model to handle long sequences while being computationally efficient. In particular, we based our code on HuggingFace’s implementation [103, 104]. Specifically, the BigBird configuration used in our model includes a block size of 64 tokens, with custom modifications to the attention mechanism. Unlike the original BigBird implementation, our configuration does not utilise window attention; each token attends only to itself and not its neighbours (in the sequence space) or tokens within the same block unless that block is selected through random attention. Additionally, global attention is restricted to the global tokens within the first block and does not extend to all tokens in the block. The random attention spans 4 blocks, a value chosen specifically for this setup. These decisions in the attention mechanism arise from the fact that we are not working with true sequential data; the ordering of photons in the input sequence is arbitrary, and we aim to avoid the model learning spurious correlations or patterns from this ordering. Our transformer encoder architecture comprises 12 attention layers, a hidden size of 384 units, GELU activations [105], and 12 attention heads. Shuffling the photon sequence mitigates the risk of the network learning undesirable block patterns.
@@
We use CD instead of the Earth Mover’s Distance (EMD) [106] due to EMD’s higher computational complexity and limitations in preserving the fidelity of detailed structures [107]. This approach facilitates learning the overall shape of the neutrino interaction rather than precisely backtracking each individual photon ray. A known limitation of CD is that it does not account for local variations in point density. In regions with low ground-truth density, a single true point may serve as the nearest neighbour for many predicted points, leading to an over-representation of these sparse areas in the loss, while regions with denser ground truth are under-represented. To overcome this limitation, the Density-Aware Chamfer Distance (DCD) [107] is adopted. The DCD augments the CD loss by incorporating a density-aware weighting scheme along with exponential mapping.
@@
For small distances, the contribution is approximately linear; for larger distances, the contribution saturates, thereby reducing the influence of outliers. The total loss applied to the transformer encoder layers comprises both the DCD and a weighted MSE loss, ensuring a balanced learning process. Deep supervision [108, 109] is applied by computing the loss at each encoder layer with exponentially increasing weights. For a transformer with encoder layers, the total loss is
@@
The model was implemented in Python 3.12.2 using PyTorch 2.5.1 [110] and PyTorch Lightning 2.4.0 [111], and trained on a single NVIDIA H100 GPU with 94 GB of memory. The training was conducted over 200 epochs with a batch size or 64 and a maximum input sequence length capped at 1024, as previously specified. The optimiser employed was AdamW [112, 113], with a learning rate initialised at , , , and a weight decay of . The learning rate schedule consisted of a single warm-up epoch, followed by cosine annealing over the remaining 199 epochs, decaying smoothly to zero. Hyperparameter optimisation was conducted using the Optuna framework, enabling efficient search space exploration through automated sampling and pruning strategies [114]. No dropout was applied, as empirical evidence indicated a detrimental effect on the regression of precise continuous values. The generalisation capacity typically provided by dropout was instead achieved through the stochastic sampling of photons per event. Given the large number of possible photon combinations, the likelihood of the model encountering the same subset of inputs across epochs is negligible.

### v4

@@ 4.1 Optical model and simulation
The interaction and propagation of particles in the scintillator volume were simulated with the Geant4 toolkit [Allison_2006_Geant4_developments_and, Agostinelli_2003_Geant4_a_simulation_toolkit, Allison_2016_Recent_developments_in]. The parameters of polyvinyltoluene-based plastic scintillator are simulated with light yield (8,700 photons / MeV), emission and absorption spectra, attenuation length (2.5 m), refractive index (1.58), and decay time ( ns) characteristic of EJ-262 [datasheet_EJ_260_262], the scintillator used for the PLATON-prototype. It is worth mentioning that the purpose of this simulation study is general and is applicable to a wide variety of organic scintillators, including those in the form of liquid, often used in giant neutrino experiments [JUNO:2021vlw, DAYABAY]. Their typical light yield is around 10,000 photons / MeV, attenuation lengths equal or above 20 m, as well as very high purity, thus, a very uniform optical performance.
Neutrinos are simulated as follows. The neutrino energy was simulated according to the flux made public by the T2K experiment, where it is used in the neutrino oscillation data analyses [Abe_2015_Measurements_of_neutrino, FluxPrediction2016]. The flux is peaked at 0.6 GeV and its broad band ranges mainly from 0.2 to 1.5 GeV, with a long tail up to 30 GeV. The T2K neutrino angular distribution is wide, with a standard deviation of about 4 meters at the near detector location [T2K:2012bge]. Thus, for simplicity we decided to simulate neutrinos on a fixed direction, parallel to both sensor planes, without loss of generality. The NEUT 5.5.0 neutrino event generator [Hayato_2021_The_NEUT_neutrino], used at the T2K experiment for neutrino data analyses, is used to simulate the neutrino interactions in the scintillator. It provides the list of final-state particles (particle types and corresponding momentum vectors) and the 3D position of the neutrino interaction vertex. Then, the propagation of the final-state particles in matter is simulated using the Geant4 toolkit [Allison_2006_Geant4_developments_and, Agostinelli_2003_Geant4_a_simulation_toolkit, Allison_2016_Recent_developments_in]. We assume the distance between the position of the ionising particle and the final emission of the scintillation light to be small compared to the 3D spatial resolution of the PLATON camera [PDG, Kolanoski:2020ksk] and, thus, no additional smearing on the position of the scintillation light emission is applied. The optical photons produced by the scintillation process are propagated using native Geant4 optical photon processes, including absorption and reflections following the Fresnel equations. The speed of photons in scintillator is about 20 cm / ns. Once the photons leave the scintillator volume, their positions and directions are stored for subsequent processing steps.
The second part of the optical simulation is based on a custom library that simulates lenses, sensors, and refractive surfaces outside the scintillator volume. It performs the ray tracing of each individual photon from the exit point on the scintillator surface to the photosensor, through the plenoptic system. Thin and thick lenses, together with their apertures, are implemented in the paraxial approximation, assuming 100% lens transparency and neglecting optical aberrations. The propagation time of each photon is computed assuming straight-line travel between optical elements and using the speed of light in vacuum. Further, general surfaces can be added to the optical stack by defining their shape using tessellated triangles in an STL file.
@@
The simulation of the SPAD array sensors was based on the parameters of SwissSPAD2 [Ulku_2019_A_512_x]. The pixel active and dead areas of the SPAD array were simulated as designed, thus with a geometrical fill factor. The photon detection efficiency (PDP) was simulated as a binomial efficiency following the corresponding wavelength spectrum. The DCR was also cross-checked against measurements performed in our laboratory, and where explicitly mentioned, extrapolated to a future, realistic performance. The timestamp of each pixel can be simulated in either self-trigger or gate mode.
The positions and the relative distances of the optical objects were adjusted to reflect the configuration of the PLATON-prototype (see Sec. 4.2 and 4.3) or that of an optimised PLATON detector (see Sec. 2.3). Although this model appears rather simple, it accurately reproduces the actual resolution of the PLATON-prototype.
@@ 4.2 Image post-processing method
Post-processing algorithms for plenoptic systems are based on parallax detection and require pattern matching between images observed by different micro-lenses [konz2016depthest]. In this reconstruction regime, a single point source would create a sub-image in a set of neighbouring micro-lenses. If the point source sits in the focus region of the plenoptic camera, each microlens images a single pixel, while a point source out of focus would create a blurred image in each participating microlens. The number of active microlenses is known as Virtual Depth (VD) and is proportional to the physical distance of the object from the main lens, allowing for depth estimation. However, we decided not to adopt this technique as it cannot be easily applied to photon-starved frames, as it would lead to the pattern matching failing. Thus, we require an optical model (see Sec. 4.1) to project each detected photon through the optics into the object space. Post-processing, based on that of a pinhole camera, is also used for plenoptic cameras in light-intensity imaging [chief-ray-hahne]. In this work, we adopt the same principle to the case of photon-starved images and, thus, developed the following method: the ray trace of every photon that was detected by a pixel of the sensor (a count) is forced to pass through the centre of the closest micro-lens, and interpolated in the object space. The optical model described in Sec. 4.1 was used for the propagation of the ray tracing: the MLA was parametrised as an array of pinholes, and each photon was assumed to have impinged at the centre of the pixel. Outlier ray traces, mainly generated by noise, but also by the adopted approximation, were iteratively rejected until either a predefined minimum number of rays with the relative closest distance was obtained or the distance between each leftover ray was below a certain threshold. The interpolation with the remaining ray traces was used to estimate the position of the light source.
@@
4.3 Calibration of the PLATON-prototype
The post-processing of the 2D images relies on the accurate calibration of the optical model of the prototype, that is, the data-driven inference of the model parameters discussed in Sec. 4.1. First, the lateral position of each micro-lens is obtained by illuminating the PLATON-prototype without the main lens with a source of parallel light rays. This was achieved using a low-power laser mounted on a custom-made motorised stage, which could be translated parallel to the MLA plane. A single greyscale intensity image was obtained from 4’096 frames and used to determine the centre of each micro-lens. The mean distance between the centre of two adjacent micro-lenses was computed to derive the parameter . The position of the micro-lenses could not be determined for the entire area of the MLA, as the glue partially covers that region (see Fig. 1). Consequently, that part of the image was not used in the post-processing. This issue reduced the field of view of the PLATON-prototype. Imaging a parallel-light source is not sufficient to infer the remaining parameters of the optical model. Hence, we acquired a dataset consisting of 12-bit images of a back light-illuminated 50 m pinhole at 168 different positions within an imaging volume of . Each image was obtained by stacking 4,096 frames, each recorded with a 10 s gate. The motorised stage, by design, allowed the light source to move in steps of . Measurements where the light source fell outside the camera’s field of view (FoV), reduced by the glue, were removed from the dataset. In total, 160 images were used to fit the parameters of the optical model using a custom genetic algorithm, in which the post-processing method was iteratively applied and the distance between the reconstructed and true positions was minimised.
@@ 4.3 Calibration of the PLATON-prototype
The optical simulation (see Sec. 4.1) was used to reproduce images from a 100 point source at the same 168 positions as those of the acquired dataset. The focal length of the micro-lens array, , was fixed to its nominal value in the datasheet of the MLA. Then, the simulated images were compared to the corresponding data samples, and good agreement was observed, indicating that the model provides a reasonable representation of the camera’s optical stack.
@@ 4.4 Data-driven likelihood reconstruction analysis
Alternative post-processing methods were also tested. The first one consists of a data-driven likelihood analysis. The intensity greyscale images for each of the 160 points scanned during the calibration campaign were used as probability density functions (PDFs). To each PDF, the distribution of the dark counts across the pixels, obtained with an independent measurement, is subtracted. This way we could estimate the intrinsic spatial resolution of the plenoptic system for the case of a negligible contamination from dark counts. This study aims to emulate the scenario of a future SPAD array chip capable of rejecting dark counts by time coincidence by constructing sampled data frames: after subtracting the independently measured dark-count map, we generate low-photon-count binary frames by sampling from the resulting intensity PDFs. See more details in the discussion in Sec. 3. ‘Sampled” data frames were obtained by randomly throwing the position of the counts from the PDF of a given position of the light source.
@@
where the first term contributes only when the given pixel in the sampled frame takes the value of 1, while the second term contributes when the pixel is 0 and weighs the probability of the pixel to be off. This likelihood definition allows for taking into account both activated and non-activated pixels. The prefactor normalises the to an average contribution per pixel, keeping its scale independent of the number of pixels entering the comparison and without affecting the position of the minimum with respect to . For each one, the was computed with respect to the PDF at every scanned position , and the position yielding the lowest value was selected as the best-fit estimate. We refer to this approach as the “likelihood” method.
Since the testable resolution of the likelihood method depends on the distance between two adjacent scanned positions, the calibration sample was augmented within a smaller portion of the volume, which was sampled with a higher density of points. At the depths of 245 mm and 345 mm from the sensor, 101 16-bit images were taken respectively, with a spacing of 0.1 mm. These were used to determine the lateral resolution. For the depth resolution, data points with a spacing of 1 mm along the optical axis, between a 230 mm and 450 mm distance from the sensor, were gathered. Each of these intensity images was used as an empirical PDF to infer the position of the point light source, and the corresponding sampled low-photon-count datasets were generated as described above.
@@ 4.4.1 Point-light spatial resolution and comparison with classical cameras
Before explicitly studying the capability of a PLATON-10cm module to detect GeV neutrinos, a simulation study was performed to evaluate the spatial resolution to a point light source corresponding to 1 MeV energy deposition. The post-processing method used to evaluate the spatial resolution for a point light source is the same as described in Sec. 4.2. For a single-point scintillation light source, creating 10,000 photons, the 3D spatial resolution is 0.3 mm. If the number of cameras is reduced to two, the 3D spatial resolution becomes 0.8 mm.
For consistency, we compared the performance of the plenoptic cameras with classical cameras, obtained by removing the MLA and adapting the lens-to-sensor distance. The post-processing adopted for the classical cameras is the same as that used for plenoptic cameras, with the difference that now the pinhole is the main lens. While a single classical camera cannot precisely reconstruct the 3D position of the light source, the resolution of two orthogonal cameras is 2.4 mm, which is three times worse than that of the corresponding plenoptic system. Moreover, we note that two classical cameras fail to reconstruct the position of the light source for approximately 23% of the events, whereas all events are successfully reconstructed for two plenoptic cameras. The resolution for eight cameras, split into two orthogonal views, is approximately four times worse than that of the PLATON-10cm detector module, which utilises eight plenoptic cameras.
@@ 4.5 Neural-network based reconstruction
The dataset used for the neural network studies comprised 1M CC-inclusive muon neutrino interactions generated as described in Sec. 2.3.2. The simulation of the detector setup is discussed in Sec. 2.3. Reconstructing 3D neutrino interactions within this setup is inherently challenging. SPAD arrays are arranged orthogonally on two sides of the scintillator block. Each interaction produces a photon-starved pattern, typically comprising between a few tens and several thousand detections across the SPADs. The task is to infer the spatial origin and topology of the interaction using only these sparse observations.
@@
To achieve this, we opt for a transformer-encoder-based architecture to fully exploit the spatial correlations among detected photons through the transformer’s attention mechanism [NIPS2017_3f5ee243, NIU202148]. Unlike traditional architectures that rely on local connectivity or fixed sequence order, the transformer’s attention mechanism enables dynamic interactions between all elements in the input. This is particularly beneficial for our application, where the input consists of sparse and spatially-distributed photon detections and where meaningful correlations may exist between distant regions of the detector. Compared to convolutional neural networks (CNNs) [firstCNN, 9451544] or graph neural nets (GNNs) [4700287, 9046288, ZHOU202057], which are constrained by local or explicitly defined receptive fields, transformers are not limited to such neighbourhoods and can integrate global spatial context more effectively. Recurrent neural networks (RNNs) [JORDAN1997471, 10.5555/553011, SHERSTINSKY2020132306] are also ill-suited for this task, as they assume a consistent sequential ordering, which does not naturally apply to unordered 2D photon hits [Xiao2021]. In contrast, the transformer’s attention mechanism provides the flexibility to learn arbitrary spatial dependencies, which is essential for resolving the complex and often ambiguous photon distributions resulting from neutrino interactions [ISLAM2024122666, 10.1145/3505244]. An alternative class of models worth considering is Neural Radiance Fields (NeRFs) [mildenhall2020nerfrepresentingscenesneural], which learn continuous volumetric scene representations from dense multi-view RGB data using ray-based modelling. While NeRFs have proven effective for reconstructing 3D geometry from 2D projections, their direct application to our setting is limited. SPAD sensors produce extremely sparse, binary detections, lacking the dense photometric supervision and calibrated viewing geometries that NeRFs rely on. Furthermore, NeRFs are primarily designed for photorealistic view synthesis, whereas our goal is to recover the underlying spatial topology of particle interactions from non-image-based, unordered data. As such, although NeRF-inspired techniques may hold promise for future work [yu2021pixelnerfneuralradiancefields, niemeyer2021regnerfregularizingneuralradiance], they would require substantial re-engineering to address the unique challenges of our domain. Moreover, more traditional reconstruction methods are discarded due to the multi-dimensional complexity of the problem and the large number of sparse photons involved. These methods often lack the flexibility and computational efficiency required to process and interpret the intricate patterns of photon detection necessary for accurate 3D reconstruction.
@@
We employ a transformer encoder to tackle this reconstruction problem (see Figure 9a). In our setup, each photon is treated as a token, embedding its 2D position on the SPAD array, the identifier of the micro-lens it passed through, and the corresponding SPAD sensor ID. Additionally, the input sequence includes a dedicated token encoding the number of sampled photons in the event and one token for each muon exiting the scintillator block (typically one), which contains the 3D exit point and direction vector. The maximum sequence length is fixed at 1,024 tokens. During training, we randomly select of photons per event (with being the number of photons in the event, further capped by the maximum sequence length) and fill the remainder of the sequence with a padding token, which is ignored during training. This stochastic subsampling means that, from the model’s perspective, the same physical event is seen with a different subset of photons at each epoch, and thus acting as an input-level regularisation/data-augmentation mechanism, reducing overfitting to specific photon patterns and improving generalisation to unseen events. Given that each token already includes spatial information, explicit positional encodings are unnecessary. Early experiments showed that the network struggled to reconstruct events when the number of detected photons was significantly lower than the maximum sequence length. We hypothesise that this limitation arises from the model’s lack of awareness of the photon count in each event. To address this, we include the aforementioned dedicated token encoding the sequence size, which resolves this issue by enabling the model to condition the number of observed photons.
@@
The quadratic complexity of the standard transformer’s attention mechanism poses computational challenges for processing events with a large number of photons. To overcome this, we use a custom-adapted version of BigBird’s sparse attention [zaheer2021bigbirdtransformerslonger], which combines global, window, and random attention within blocks of tokens to achieve linear complexity while effectively modelling long-range dependencies. Global attention allows specific tokens to attend to all others, window attention limits tokens to attend only to their neighbours, and random attention randomly connects tokens. This hybrid approach allows the model to handle long sequences while being computationally efficient. In particular, we based our code on HuggingFace’s implementation [wolf2020huggingfacestransformersstateoftheartnatural, bigbird-huggingface]. Specifically, the BigBird configuration used in our model includes a block size of 64 tokens, with custom modifications to the attention mechanism. Unlike the original BigBird implementation, our configuration does not utilise window attention; each token attends only to itself and not its neighbours (in the sequence space) or tokens within the same block unless that block is selected through random attention. Additionally, global attention is restricted to the global tokens within the first block and does not extend to all tokens in the block. The random attention spans 4 blocks, a value chosen specifically for this setup. These decisions in the attention mechanism arise from the fact that we are not working with true sequential data; the ordering of photons in the input sequence is arbitrary, and we aim to avoid the model learning spurious correlations or patterns from this ordering. Our transformer encoder architecture comprises 12 attention layers, a hidden size of 384 units, GELU activations [hendrycks2023gaussianerrorlinearunits], and 12 attention heads. Shuffling the photon sequence mitigates the risk of the network learning undesirable block patterns.
@@
We use CD instead of the Earth Mover’s Distance (EMD) [710701] due to EMD’s higher computational complexity and limitations in preserving the fidelity of detailed structures [wu2021densityawarechamferdistancecomprehensive]. This approach facilitates learning the overall shape of the neutrino interaction rather than precisely backtracking each individual photon ray. A known limitation of CD is that it does not account for local variations in point density. In regions with low ground-truth density, a single true point may serve as the nearest neighbour for many predicted points, leading to an over-representation of these sparse areas in the loss, while regions with denser ground truth are under-represented. To overcome this limitation, the Density-Aware Chamfer Distance (DCD) [wu2021densityawarechamferdistancecomprehensive] is adopted. The DCD augments the CD loss by incorporating a density-aware weighting scheme along with exponential mapping.
@@
For small distances, the contribution is approximately linear; for larger distances, the contribution saturates, thereby reducing the influence of outliers. The total loss applied to the transformer encoder layers comprises both the DCD and a weighted MSE loss, ensuring a balanced learning process. Deep supervision [lee2014deeplysupervisednets, li2022comprehensivereviewdeepsupervision] is applied by computing the loss at each encoder layer with exponentially increasing weights. For a transformer with encoder layers, the total loss is
@@
The model was implemented in Python 3.12.2 using PyTorch 2.5.1 [NEURIPS2019_9015] and PyTorch Lightning 2.4.0 [Falcon_PyTorch_Lightning_2019], and trained on a single NVIDIA H100 GPU with 94 GB of memory. The training was conducted over 200 epochs with a batch size or 64 and a maximum input sequence length capped at 1024, as previously specified. The optimiser employed was AdamW [Kingma-2014-adam, loshchilov2019decoupledweightdecayregularization], with a learning rate initialised at , , , and a weight decay of . The learning rate schedule consisted of a single warm-up epoch, followed by cosine annealing over the remaining 199 epochs, decaying smoothly to zero. Hyperparameter optimisation was conducted using the Optuna framework, enabling efficient search space exploration through automated sampling and pruning strategies [optuna_2019]. No dropout was applied, as empirical evidence indicated a detrimental effect on the regression of precise continuous values. The generalisation capacity typically provided by dropout was instead achieved through the stochastic sampling of photons per event. Given the large number of possible photon combinations, the likelihood of the model encountering the same subset of inputs across epochs is negligible. All results presented in Sec. 2.3.2 were obtained on the held-out test set from a train/validation/test split, respectively.